OpenStudy (alivejeremy):

Numbers p and q are negative integers. Which statements are always true? Choose all answers that are correct. A. p + q is a negative integer B. p • q is a positive integer C. p – q is a negative integer D. is a negative integer

OpenStudy (anonymous):

need hlp

OpenStudy (alivejeremy):

oh god

OpenStudy (anonymous):

no

OpenStudy (alivejeremy):

yea

OpenStudy (anonymous):

no

OpenStudy (alivejeremy):

help plz

OpenStudy (anonymous):

hi

OpenStudy (alivejeremy):

hey

OpenStudy (alivejeremy):

can u help me

OpenStudy (anonymous):

what grade r u in

OpenStudy (anonymous):

looks like 7th

OpenStudy (anonymous):

its b

OpenStudy (alivejeremy):

its more than one

OpenStudy (anonymous):

it is not b

OpenStudy (anonymous):

what grade r u in

OpenStudy (anonymous):

yes it is

OpenStudy (alivejeremy):

7th

OpenStudy (alivejeremy):

wbu

OpenStudy (anonymous):

do u go 2 k12

OpenStudy (alivejeremy):

yea

OpenStudy (anonymous):

p + q is a negative integer p • q is a positive integer

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

i am in k12 7th and i have 2 more lessons to finish math

OpenStudy (anonymous):

medal?????

OpenStudy (anonymous):

p • q is a positive integer

OpenStudy (anonymous):



OpenStudy (anonymous):

@siblings is in 6th grade

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

6th

OpenStudy (alivejeremy):

lol

OpenStudy (anonymous):

complete the answer 4 d

OpenStudy (anonymous):

so what there is no problem with my grade

OpenStudy (anonymous):

a is one answer

OpenStudy (alivejeremy):

i think its A,B,D

OpenStudy (anonymous):

a and b is you answer

OpenStudy (anonymous):

b no

OpenStudy (anonymous):

d is not full

OpenStudy (anonymous):



OpenStudy (anonymous):

d isX

OpenStudy (alivejeremy):

A and C

OpenStudy (anonymous):

A and B

OpenStudy (anonymous):

c no

OpenStudy (alivejeremy):

Is the set of multiples of 4 closed under addition? Explain why or provide a counterexample if not. A. Yes, because the sum of any two multiples of 4 is also a multiple of 4. B. Yes, because the sum of any two multiples of 4 is also a multiple of 8. C. No, and a counterexample is –4 + (–8) = –12. D. No, and a counterexample is 26 + 16 = 42.

OpenStudy (anonymous):

d is wrong

OpenStudy (anonymous):

a and d r going 2 be divided

OpenStudy (anonymous):

a and d

OpenStudy (alivejeremy):

Is the set of multiples of 4 closed under addition? Explain why or provide a counterexample if not. A. Yes, because the sum of any two multiples of 4 is also a multiple of 4. B. Yes, because the sum of any two multiples of 4 is also a multiple of 8. C. No, and a counterexample is –4 + (–8) = –12. D. No, and a counterexample is 26 + 16 = 42.

OpenStudy (anonymous):

99.99% sure

OpenStudy (anonymous):

a and d

OpenStudy (alivejeremy):

lol 100%

OpenStudy (anonymous):

100%

OpenStudy (alivejeremy):

only one for this one

OpenStudy (anonymous):

100000000000000000000%